According to the rocket scientists, what angle doe
Answer #1

let, tan(45) = y/L


use length contraction equation, L' = L*sqrt(1 - (v/c)^2)

= L*sqrt(1 - (1*10^6/(3*10^8))^2)

= L*0.999994


let theta is the angle made by the E with rocket with respect rocket.

so, tan(theta) = y/L'

= y/(L*0.999994)

= tan(45)/0.999994

= 1.000006

theta = tan^-1(1.000006)

= 45.00017 degrees